OpenStudy (anonymous):

solve for x. 4(x-9)-6=-6(-3x+6)-6x

OpenStudy (anonymous):

First distribute the 4 and -6:\[4x - 36 - 6 = 18x - 36 - 6x\]Then combine like terms:\[4x - 42 = 12x - 36\]Then use algebra to rearrange the equation so that all terms with x are on one side and everything else is on the other:\[-6 = 8x\]Then divide by 8 to get x by itself:\[-\frac{3}{4} = x\]

OpenStudy (anonymous):

thank u! ur a smart kid i guess